larry mitchell invested part of his 14,000 advance at 7% annual simple interest and the rest at 3% annual simple interest. if his total yearly interest from both accounts was $740 find the amount invested at each rate?
Answer by Boreal(15235)

x=amount at 7%
14000-x =amount at 3%
.07x+(0.03)*(14000-x)=740
.07x+420-.03x=740
.04x=320
x=$8000 at 7%=$560
14000-x=$6000 at 3%=$180
They add to $740